# Unleashing the Potential of Tracklets for Unsupervised Video Person Re-Identification

### v1

@@ III-B Noise-Filtered Tracklet Partition
Performing clustering on tracklet features can generate inaccurate pseudo labels since there are numerous harmful frames and unexpected variations within tracklets. Thus, we propose the Noise-Filtered Tracklet Partition (NFTP) module to eliminate these negative impacts, further promoting our proposed self-supervised refined clustering.
@@
Noisy frame filtering. At the beginning of each epoch, we use the latest model to extract the frame features $\left\{f_{i,j}\right\}_{j=1}^{L_{i}}$ of each tracklet $T_{i}$, and then average them to get the center feature $C_{i}$ as follows:
$C_{i}=\frac{1}{L_{i}}\sum_{j=1}^{L_{i}}f_{i,j}$ (6)
@@
$q_{i}=\frac{1}{L_{i}*\delta}\sum_{j=1}^{L_{i}}{\rm dist}\left(f_{i,j},C_{i}\right)$ (7)
where ${\rm dist}\left(f_{i,j},C_{i}\right)$ is computed as ${\left(1-f_{i,j}\cdot C_{i}\right)}^{2}$ and $f_{i,j}\cdot C_{i}$ denotes the cosine similarity between each frame $f_{i,j}$ and the center feature $C_{i}$. $\delta$ is a hyperparameter that adjusts the degree of noise filtering. For each frame $x_{i,j}$, if ${\rm dist}\left(f_{i,j},C_{i}\right)>q_{i}$, it is considered a noisy frame. The threshold $q_{i}$ is dynamically adjusted according to the frame features and the length of tracklets, so that noises within tracklets can be effectively filtered. To help understand the effectiveness of noisy frame filtering, we provide more discussions on the impact of $\delta$ on the adaptive threshold and the visualization of noisy frames in Figs. 5 and 7, respectively.
Tracklet partition. To alleviate unexpected variations within tracklets, given each noise-filtered tracklet $\hat{T}_{i}=\left\{x_{i,j}\right\}_{j=1}^{\hat{L}_{i}}$, where $\hat{L}_{i}$ is the number of frames, we employ a sequentially equidistant partition strategy to divide each tracklet $\hat{T}_{i}$ into multiple “sub-tracklets”. Formally,
$\hat{T}_{i}=\left\{\hat{ST}_{i,t}\right\}_{t=1}^{n^{\prime}_{i}},~{}\hat{ST}_{i,t}=\{x_{i,j}\}^{t\cdot l}_{j=1+(t-1)l}$ (8)
@@
III-C Self-Supervised Refined Clustering
@@
Reliable sub-cluster generation. DBSCAN [45] is employed for clustering, which utilizes a scanning radius hyper-parameter, denoted as $eps$, to regulate the size of the cluster. To ensure the consistency of identities within clusters, we set a strict constraint condition of small $eps$. In this way, only more reliable sub-tracklets, i.e., sub-tracklets with relatively close feature distance with each other, are considered to share the same identity. We can get the labeled dataset $\hat{X}=\{(\hat{ST}_{i},\hat{y_{i}})\}_{i=1}^{N^{s}}$, where $\hat{y_{i}}$ denotes the pseudo label of the $i$-$th$ sub-tracklet. Since our proposed NFTP module has filtered noises within sub-tracklets, the distance between sub-tracklets of different identities is extended. They are harder to be clustered under the strict clustering constraint, alleviating the problem in Fig. 1 (a). However, there are still spatial-temporal changes in each tracklet, such as complicated backgrounds or illumination variations, so sub-tracklets of the same tracklet can also be grouped into different clusters under such a strict clustering process. Therefore, we further merge the obtained clusters, which are viewed as “sub-clusters”, to form the ultimate clustering results.
Sub-cluster merging. The natural self-supervised signal from the tracklet partition in Sec. III-B can serve as a guide for sub-cluster merging. Since sub-tracklets are obtained through the sequential partition, various sub-tracklets from the same tracklet indicate the same identity. Additionally, as observed in Fig. 1, cross-camera tracklets may contain similar local areas, so that cross-camera sub-tracklets of the same identity can be clustered into the same sub-cluster. With the merging process based on the self-supervised signal from the tracklet partition, various cross-camera tracklets of the same identity can be effectively aggregated together.
Figure 3: Illustration of our proposed progressive strategy for sub-cluster merging. Based on the reliable sub-clusters, only directly reachable sub-clusters are merged during the early stages of training, while all reachable sub-clusters are merged later.
Progressive merging strategy. Considering the feature extraction capability of the model is inferior at the early training stages, and improves as training progresses, we incorporate a progressive strategy to facilitate effective sub-cluster merging as shown in Fig. 3. For a clear description, we define two terms here:
@@
During the early stages of training, due to the poor feature extraction ability, sub-tracklets from different identities may be assigned to the same sub-cluster. To avoid error propagation during the merging step, we only merge directly reachable sub-clusters (Def. 1) at the early training stages. As training goes on, the model becomes more robust and sub-clusters become more reliable. Therefore, we propose to merge all reachable sub-clusters (Def. 2) after the learning rate decay. Benefiting from the progressive merging strategy, reliable clusters can effectively promote feature learning of the model, making it less affected by clustering noises.
After sub-cluster merging, each cluster with refined pseudo label $\tilde{y_{i}}$ is composed of multiple sub-clusters. As illustrated in Fig. 3, we denote the set of all pseudo labels $\hat{y}$ of these sub-clusters as $P_{i}$. Finally, with our proposed self-supervised refined clustering, we can get the dataset $\tilde{X}=\{(\hat{ST}_{i},\hat{y_{i}},\tilde{y_{i}})\}_{i=1}^{N^{s}}$ with two kinds of pseudo labels for each sub-tracklet.
III-D Training
Generally, we follow the training framework as introduced in Sec. III-A, except that we apply our proposed SSR-C at the sub-tracklet level. The training procedure of our proposed SSR-C framework is provided in Alg. 1. At the beginning of each epoch, we first freeze the model and apply our proposed NFTP module to obtain noisy-filtered sub-tracklets. Then, we extract the features of noisy-filtered sub-tracklets via Eq. 1, with the only difference that we randomly sample $M$ frames for each noisy-filtered sub-tracklet rather than the original tracklet. Subsequently, our proposed self-supervised refined clustering method is conducted based on the sub-tracklet features. After that, the memories $\boldsymbol{V}$ and $\boldsymbol{V^{h}}$ are initialized with sub-cluster feature centers. With abundant pseudo labels of sub-tracklets and feature centers generated by self-supervised refined clustering, we design an effective loss, namely Class Smoothing Classification (CSC) loss, to effectively guide the model learning at the fine-grained sub-tracklet level, while both $\boldsymbol{V}$ and $\boldsymbol{V^{h}}$ are updated using sub-tracklet features with Eq. 3. By iteratively performing our proposed self-supervised refined clustering with NFTP and effective model optimization, the feature extraction capability and robustness of the model can be significantly enhanced.
Input: Unlabeled training set $X=\left\{T_{1},T_{2},\cdots,T_{N}\right\}$; Number of tracklets $N$; CNN model $\phi\left(\theta\right)$; Total training epoch $EP$; Training iterations per epoch $ITER$.
1 for epoch $\leftarrow$ 1 to $EP$ do
@@ 2       Freeze the model $\phi\left(\theta\right)$;
3       // Perform the NFTP module
4       for $i$leftarrow 1 to $N$ do
5             $C_{i}$leftarrow Compute the center feature of $T_{i}$ via Eq. 6;
6             $\hat{T}_{i}$leftarrow Filter noisy frames in $T_{i}$ via Eq. 7;
7             $\hat{ST}$leftarrow Partition $\hat{T}_{i}$ into sub-tracklets via Eq. 8;
8
@@
10      $X\leftarrow\{\hat{ST}_{1},\hat{ST}_{2},\cdots,\hat{ST}_{N^{s}}\}$
11       // Perform self-supervised refined clustering
12       for $i$leftarrow 1 to $N^{s}$ do
@@ 19       Initialize $\boldsymbol{V}$ and $\boldsymbol{V}^{h}$ with sub-cluster feature centers;
20       // Train the model with $\tilde{X}$
@@ 21       Unfreeze the model $\phi\left(\theta\right)$;
22       for iter $\leftarrow$ 1 to $ITER$ do
23             $\{\hat{ST}_{i}\}_{i=1}^{B}$leftarrow Sample a training batch;
24             for $i$leftarrow 1 to $B$ do
25                   $v_{i}\leftarrow$ Extract the feature of $\hat{ST}_{i}$ via Eq. 1;
26                   Optimize $\theta$ via Eqs. 9, 10 and 5;
27
@@ 23             $\{\hat{ST}_{i}\}_{i=1}^{B}$leftarrow Sample a training batch;
29            Update $\boldsymbol{V}$ and $\boldsymbol{V}^{h}$ via Eq. 3;
@@
Class smoothing classification loss. Based on the abundant pseudo labels $\hat{y}$ and $\tilde{y}$ as well as the $P_{i}$ set obtained with our proposed self-supervised refined clustering, we define the Class Smoothing Classification (CSC) loss as follows:
@@
$\mathcal{L}_{c}^{csc}\!=-\!\!\!~{}\sum_{j\in P_{i}}s_{j}\log\frac{\exp\left(\boldsymbol{v_{i}}^{\top}\!\cdot\!\boldsymbol{V}_{j}/\tau\right)}{\exp\left(\boldsymbol{v_{i}}^{\top}\!\cdot\!\boldsymbol{V}_{j}/\tau\right)\!+\!\!\!\!\sum\limits_{k\notin P_{i}}\exp\left(\boldsymbol{v_{i}}^{\top}\!\cdot\!\boldsymbol{V}_{k}/\tau\right)}$ (9)
@@
$s_{j}=\left\{\begin{array}[]{lr}1-\lambda+\frac{\lambda}{K},&j=\hat{y}_{i}\\ \frac{\lambda}{K},&{\rm otherwise}\end{array}\right.,\quad\forall j\in P_{i}$ (10)
where $\lambda$ is the smoothing term, $K$ is the cardinal number of the positive set $P_{i}$. Similarly, we can define $\mathcal{L}_{h}^{csc}$ and apply the overall loss functions as Eq. 5.
@@
We note that some works [25, 37, 49, 50] simply select $K$ potential positive samples to mitigate the impact of clustering errors. They adopt feature similarity as metrics and treat $K$ as a predefined hyper-parameter, which limits their methods in two ways. (1) The feature extraction ability of the model is changing at different training stages, making it difficult to determine the number of positive samples. What’s worse, there are various positive samples for different people, so a static $K$ makes lots of samples mistakenly optimized. (2) Some hard potential positive samples would not be identified simply based on feature similarity. In contrast, our method effectively addresses these problems by determining the value of $K$ dynamically based on both feature similarities and guidance of self-supervised signals. With the empowerment of our proposed NFTP module and self-supervised refined clustering, accurate $K$ positive sub-clusters can guide the learning of the model in the fine-grained sub-tracklet embedding space, significantly improving its discriminative ability.
@@
Our proposed CSC loss also exhibits significant differences from the widely-used label smoothing regularization [51]. Formally, we set small non-zero values for the weight of potential positive classes in Eq. 10, but the weights of negative classes are still 0. In contrast, label smoothing regularization sets all weights of negative classes to small non-zero values. Moreover, CSC loss aims to optimize the model in a more fine-grained sub-tracklet embedding space leveraging our refined pseudo labels $\hat{y_{i}}$ and $\tilde{y_{i}}$, while label smoothing regularization is used to avoid overfitting.
@@ IV Experiments
IV-A Datasets and Evaluation Protocol
Datasets. We evaluate our method on two widely used video person Re-ID datasets, i.e., MARS [52] and DukeMTMC-VideoReID [21]. MARS is captured by six cameras on a university campus. It contains 17,503 tracklets for 1,261 distinct identities and 3,248 distractor tracklets, which are divided into 625 identities for training and 636 identities for testing. DukeMTMC-VideoReID (Duke-V) is derived from DukeMTMC [53], which is captured in outdoor scenes with noisy backgrounds and suffers from illumination, pose, and viewpoint changes and occlusions. Duke-V is split following the protocol [54], with 2,196 tracklets of 702 identities for training and 2,636 tracklets of other 702 identities for testing.
Evaluation metrics. The Cumulative Matching Characteristic (CMC) at Rank-1/5/10, and the mean Average Precision (mAP) are used for performance evaluation.
IV-B Implementation Details
We adopt ResNet-50 [55] pre-trained on ImageNet [56] as our backbone model and use GeM pooling [57] to obtain the feature embedding of each frame. During training, we set the batch size to 32 and total training epochs to 150. Each tracklet is sequentially partitioned every $l=32$ frames to obtain sub-tracklets. For each sub-tracklet, we randomly sample $M=8$ frames with a stride of 4 to form a sub-tracklet segment. Images are resized to $256\times 128$ and augmented by horizontal flipping and random erasing [58]. As for the optimizer, AdamW with a weight decay of 0.0005 is adopted to update the parameters. The learning rate is initialized to $3.5\times 10^{-4}$ with a decay factor of 0.1 every 50 epochs. At the beginning of each epoch, we use DBSCAN [45] and Jaccard distance [59] for clustering with the $eps$ set to 0.25 and the minimal number of samples for each cluster set to 2. For hyper-parameters, $\lambda$ and $\delta$ are set to 0.1 and 0.7, respectively. Following [48], we simply set $\tau$, $\alpha$, $\gamma_{1}$, and $\gamma_{2}$ to 0.05, 0.1, 0.5, and 0.25, respectively.
[…]
Methods Extra Info. MARS Duke-V

### v2

@@ III-C Noise-Filtered Tracklet Partition
Performing clustering on tracklet features can generate inaccurate pseudo labels since there are numerous harmful frames and unexpected variations within tracklets. Thus, we propose the Noise-Filtered Tracklet Partition (NFTP) module to eliminate these negative impacts, further promoting our self-supervised refined clustering.
III-C1 Noise Filtering
@@
At the beginning of each epoch, we use the latest model to extract the frame features $\left\{f_{i,j}\right\}_{j=1}^{L_{i}}$ of each tracklet $T_{i}$, and then average them to obtain the center feature $C_{i}$ as follows:
$C_{i}=\frac{1}{L_{i}}\sum_{j=1}^{L_{i}}f_{i,j}.$ (7)
@@
$q_{i}=\frac{1}{L_{i}*\delta}\sum_{j=1}^{L_{i}}{\rm dist}\left(f_{i,j},C_{i}\right),$ (8)
where ${\rm dist}\left(f_{i,j},C_{i}\right)$ is computed as ${\left(1-f_{i,j}\cdot C_{i}\right)}^{2}$ and $f_{i,j}\cdot C_{i}$ denotes the cosine similarity between each frame $f_{i,j}$ and the center feature $C_{i}$. $\delta$ is a hyperparameter that adjusts the degree of noise filtering. For each frame $x_{i,j}$, if ${\rm dist}\left(f_{i,j},C_{i}\right)>q_{i}$, it is considered a noisy frame. The threshold $q_{i}$ is dynamically adjusted according to the frame features and the length of tracklets, so that noises within tracklets can be effectively filtered. To help understand the effectiveness of the design of dynamic noise filtering, we provide more discussions on the impact of $\delta$ on adaptive threshold and the visualization of noisy frames in Figs. 5 and 7, respectively.
III-C2 Tracklet Partition
To alleviate unexpected variations within tracklets, given each noise-filtered tracklet $\hat{T}_{i}=\left\{x_{i,j}\right\}_{j=1}^{\hat{L}_{i}}$, where $\hat{L}_{i}$ is the number of frames, we employ a sequentially equidistant partition strategy to divide each tracklet $\hat{T}_{i}$ into multiple “sub-tracklets”. Formally,
$\hat{T}_{i}=\left\{\hat{ST}_{i,t}\right\}_{t=1}^{n^{\prime}_{i}},~\hat{ST}_{i,t}=\{x_{i,j}\}^{t\cdot l}_{j=1+(t-1)l},$ (9)
@@ III-C2 Tracklet Partition
III-D Self-Supervised Refined Clustering
@@
III-D1 Reliable Sub-Cluster Generation
DBSCAN [51] is employed for clustering, which utilizes a scanning radius hyper-parameter, denoted as epseps, to regulate the size of the cluster. To ensure the consistency of identities within clusters, we set a strict constraint condition of small epseps. In this way, only more reliable sub-tracklets, i.e., sub-tracklets with relatively close feature distance with each other, are considered to share the same identity. After clustering, each cluster is assigned with a different pseudo label $\hat{y}$, and we can get the labeled dataset $\hat{X}=\{(\hat{ST}_{i},\hat{y_{i}})\}_{i=1}^{N^{s}}$, where $\hat{y_{i}}$ denotes the pseudo label of the $i$-th sub-tracklet. Since our proposed NFTP module has filtered noises within sub-tracklets, the distance between sub-tracklets of different identities is extended. They are harder to be clustered under the strict clustering constraint, alleviating the problem in Fig. 1 (a). However, there are still spatial-temporal changes in each tracklet, such as complicated backgrounds or illumination variations, so sub-tracklets of the same tracklet can also be grouped into different clusters under such a strict clustering process. Therefore, we further merge the obtained clusters, which are viewed as “sub-clusters”, to form the ultimate clustering results.
III-D2 Progressive Sub-Cluster Merging
The natural self-supervised signal from tracklet partitioning in Sec. III-C2 can serve as a guide for sub-cluster merging. Since sub-tracklets are obtained through the sequential partition, various sub-tracklets from the same tracklet indicate the same identity. Additionally, as observed in Fig. 1, cross-camera tracklets may contain similar local areas, so that cross-camera sub-tracklets of the same identity can be clustered into the same sub-cluster. With the merging process based on the self-supervised signal from the tracklet partition, various cross-camera tracklets of the same identity can be effectively aggregated together.
Figure 3: Illustration of our proposed progressive sub-cluster merging. Different shapes (e.g., triangle and circle) correspond to different tracklets, different colors (e.g., red and blue) denote different identities, and various transparencies of the same color represent the sub-tracklets of each tracklet. Based on the generated reliable sub-clusters, only directly reachable sub-clusters are merged during the early stages of training, while all reachable sub-clusters are merged later.
Considering the feature extraction capability of the model is inferior at the early training stages, and improves as training progresses, we incorporate a progressive strategy to facilitate effective sub-cluster merging as shown in Fig. 3. For a clear description, we define two terms here:
@@ III-D2 Progressive Sub-Cluster Merging
During the early stages of training, due to the poor feature extraction ability, sub-tracklets from different identities may be assigned to the same sub-cluster. To avoid error propagation during the merging step, we only merge directly reachable sub-clusters (Def. 1) at the early training stages. As training goes on, the model becomes more robust and sub-clusters become more reliable. Therefore, we propose to merge all reachable sub-clusters (Def. 2) after the learning rate decay. Benefiting from the progressive sub-cluster merging strategy, reliable clusters can effectively promote feature learning of the model, making it less affected by clustering noises.
After sub-cluster merging, each merged cluster is assigned with a different refined pseudo label (denoted by $\tilde{y}$), and all the sub-tracklets in each merged cluster share the same refined pseudo label. Finally, with our proposed self-supervised refined clustering, we can get the dataset $\tilde{X}=\{(\hat{ST}_{i},\hat{y_{i}},\tilde{y_{i}})\}_{i=1}^{N^{s}}$ with two kinds of pseudo labels for each sub-tracklet. As introduced in Sec. III-E, the abundant pseudo labels $\hat{y}$ and $\tilde{y}$ enable effective optimization in the fine-grained sub-tracklet embedding space.
Input: Unlabeled training set $X=\left\{T_{1},T_{2},\cdots,T_{N}\right\}$; Number of tracklets NN; CNN model $\phi\left(\theta\right)$; Total training epoch EPEP; Training iterations per epoch ITERITER; Training batch size $B$.
1 for epoch $\leftarrow$ 1 to EPEP do
@@ 2  Freeze the model $\phi\left(\theta\right)$;
  /* Perform the NFTP module. */
3  for $i$ $\leftarrow$ 1 to NN do
4     $C_{i}$ $\leftarrow$ Compute the center feature of $T_{i}$ via Eq. 7;
5     $\hat{T}_{i}$ $\leftarrow$ Filter noisy frames in $T_{i}$ via Eq. 8;
6     $\hat{ST} \leftarrow$ Partition $\hat{T}_{i}$ into sub-tracklets via Eq. 9;
7
@@
9 $X\leftarrow\{\hat{ST}_{1},\hat{ST}_{2},\cdots,\hat{ST}_{N^{s}}\}$;
10
 /* Perform self-supervised refined clustering. */
11  for $i$ $\leftarrow$ 1 to $N^{s}$ do
@@ 18  Initialize $\boldsymbol{V}$ and $\boldsymbol{V}^{h}$ with sub-cluster feature centers;
19
 /* Train the model with $\tilde{X}$. */
@@ 20  Unfreeze the model $\phi\left(\theta\right)$;
21  for iter $\leftarrow$ 1 to ITERITER do
22     $\{\hat{ST}_{i}\}_{i=1}^{B}$ $\leftarrow$ Sample a training batch;
23     for $i$ $\leftarrow$ 1 to $B$ do
24        $\boldsymbol{v}_{i}\leftarrow$ Extract the feature of $\hat{ST}_{i}$ via Eq. 1;
25        Optimize $\theta$ via Eqs. 10, 11 and 6;
26
@@ 22     $\{\hat{ST}_{i}\}_{i=1}^{B}$ $\leftarrow$ Sample a training batch;
28    Update $\boldsymbol{V}$ and $\boldsymbol{V}^{h}$ with sub-tracklet features via Eq. 4;
29
@@
III-E Class Smoothing Classification Loss
With abundant pseudo labels of sub-tracklets, we design an effective loss, namely Class Smoothing Classification (CSC) loss, to effectively guide the model learning at the fine-grained sub-tracklet level. As shown in Figs. 2 and 3, after adopting our self-supervised refined clustering, each merged cluster (assigned with refined pseudo label $\tilde{y}$) is composed of multiple sub-clusters (assigned with pseudo label $\hat{y}$). Therefore, each refined pseudo label $\tilde{y}_{i}$ correspond to more than one $\hat{y}$. We denote the set of all $\hat{y}$ corresponding to each $\tilde{y_{i}}$ as $P_{i}$ (as shown in Fig. 3), and define the CSC loss as follows:
@@ III-E Class Smoothing Classification Loss
$\mathcal{L}_{c}^{csc}\!=-\!\!\!~\sum_{j\in P_{i}}s_{j}\log\frac{\exp\left(\boldsymbol{v_{i}}^{\top}\!\cdot\!\boldsymbol{V}_{j}/\tau\right)}{\exp\left(\boldsymbol{v_{i}}^{\top}\!\cdot\!\boldsymbol{V}_{j}/\tau\right)\!+\!\!\!\!\sum\limits_{k\notin P_{i}}\exp\left(\boldsymbol{v_{i}}^{\top}\!\cdot\!\boldsymbol{V}_{k}/\tau\right)},$
@@
$s_{j}=\left\{\begin{array}[]{lr}1-\lambda+\frac{\lambda}{K},&j=\hat{y}_{i}\\ \frac{\lambda}{K},&{\rm otherwise}\end{array}\right.,\quad\forall j\in P_{i},$ (11)
where $\lambda$ is the smoothing term, $K$ is the cardinal number of the positive set $P_{i}$, and $\boldsymbol{V}$ is the memory, which is initialized with sub-cluster feature centers and updated using sub-tracklet features following Eq. 4. Similarly, we can define $\mathcal{L}_{h}^{csc}$ based on $\boldsymbol{V^{h}}$.
@@
We note that some works [27, 39, 54, 55] simply select $K$ potential positive samples to mitigate the impact of clustering errors. They adopt feature similarity as metrics and treat $K$ as a predefined hyper-parameter, which limits their methods in two ways. (1) The feature extraction ability of the model is changing at different training stages, making it difficult to determine the number of positive samples. There are also various positive samples for different people, so a static $K$ makes lots of samples mistakenly optimized. (2) Some hard potential positive samples would not be identified simply based on feature similarity. In contrast, our method effectively addresses these problems by determining the value of $K$ dynamically based on both feature similarities and guidance of self-supervised signals. With the empowerment of our proposed NFTP module and self-supervised refined clustering, accurate $K$ positive sub-clusters can guide the learning of the model in the fine-grained sub-tracklet embedding space, improving its discriminative ability.
@@
Our proposed CSC loss also exhibits significant differences from the widely-used label smoothing regularization [56]. Formally, we set small non-zero values for the weight term (i.e., $s_{j}$ in Eq. 11) of potential positive classes, but the weight terms of negative classes are still 0. In contrast, label smoothing regularization sets the weight terms of all negative classes to small non-zero values. Moreover, CSC loss aims to optimize the model in a more fine-grained sub-tracklet embedding space leveraging our refined pseudo labels $\hat{y_{i}}$ and $\tilde{y_{i}}$, while label smoothing regularization is used to avoid overfitting.
III-F Training Details of SSR-C
Generally, we follow the training framework as introduced in Sec. III-B, except that our proposed SSR-C optimizes the model at the sub-tracklet level. The training procedure of our proposed SSR-C framework is provided in Alg. 1. At the beginning of each epoch, we first freeze the model and apply our proposed NFTP module to obtain noisy-filtered sub-tracklets. Then, we extract the features of noisy-filtered sub-tracklets via Eq. 1, with the only difference that we randomly sample $M$ frames for each noisy-filtered sub-tracklet rather than original tracklet. Subsequently, our proposed self-supervised refined clustering method is conducted based on the sub-tracklet features. After that, the memories $\boldsymbol{V}$ and $\boldsymbol{V^{h}}$ are initialized with sub-cluster feature centers, and updated with sub-tracklet features via Eq. 4. Meanwhile, our proposed CSC loss is used to guide the model learning following Eq. 6. By iteratively performing our proposed self-supervised refined clustering with NFTP and effective model optimization, the feature extraction capability and robustness of the model can be enhanced.
@@ IV Experiments
IV-A Datasets
We evaluate our method on two widely used video person Re-ID datasets, i.e., MARS [57] and DukeMTMC-VideoReID [23]. MARS is captured by six cameras on a university campus. It contains 17,503 tracklets for 1,261 distinct identities and 3,248 distractor tracklets, which are divided into 625 identities for training and 636 identities for testing. DukeMTMC-VideoReID (Duke-V) is derived from DukeMTMC [58], which is captured in outdoor scenes with noisy backgrounds and suffers from illumination, pose, and viewpoint changes and occlusions. Duke-V is split following the protocol [59], with 2,196 tracklets of 702 identities for training and 2,636 tracklets of other 702 identities for testing.
IV-B Evaluation Metrics
The Cumulative Matching Characteristic (CMC) [60] at Rank-1/5/10, and the mean Average Precision (mAP) [61] are used for performance evaluation.
TABLE I: Comparison with state-of-the-art methods on two video person Re-ID datasets, i.e., MARS and Duke-V. “Label” denotes requiring one tracklet annotation for each person. “Cameras” denotes requiring camera labels. Supervised methods using manually annotated identity labels are represented by $\dagger$}”. The best results are shown in bold.
Methods Year Extra Info. MARS Duke-V